The arithmetc sequence has a common difference equal to 10 and its 6th term is equal to 52 find its 15th term
Answer by fractalier(6550)

If the sixth term, a6 = 52, and d = 10, then
a15 = a6 + (15 - 6)d
a15 = 52 + 9(10) = 142